# Excited State Densities from Time-Dependent Density Functional Response Theory

### v1

@@ 4 Conclusions
In this work we developed a general formalism for obtaining the real-space excited-state densities from linear response TDDFT using time-independent perturbation theory. While the approach is equivalent to earlier work 4, 5 presented using the stationary property of excitation energies and the variational principle, our work here goes beyond the restriction to the adiabatic approximation assumed there. We observe that even with two contrasting approximations (LDA and EXX), we obtain accurate results for the lowest local and charge-transfer excitations between closed-shell fragments. Within the small-matrix approximation, which in its essence retains a single KS transition in its calculation of the energy, the excited-state density still incorporates corrections from all possible KS transitions, leading to a significant improvement over the KS approximation. Our results show the importance of an accurate KS potential in the calculation of excited-state density, which is better provided by adiabatic EXX than the LDA. For the future calculations of excited-state densities in real systems, our results are consistent with previous works 8, 9, 10, 11, 12, 13, 14 that suggest the use of hybrid functionals with large fraction of exact exchange as a more reliable choice.
The practical use of the general formalism is limited to the adiabatic approximation and to proposed approximations that are available for frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ kernels. (A general-purpose frequency-dependent kernel does not exist). We show that for doubly-excited states, the densities from the DSPA and DSMA-based formulation, both previously shown to be good for excitation energies and DSMA for oscillator strengths, provides accurate double excitation densities, which is not achievable through adiabatic approximation.
Benchmarking on a wider set of real molecules is required to further probe the efficiency and accuracy of the linear-response formalism developed for real-space excited-state densities for states of both single and double excitation character, as has been explored for dipole moments restricted to the adiabatic approximation (single excitations) in earlier works 8, 9, 10, 11, 12, 13, 14, 15. Beyond static properties, the formalism can be used in the response-reformulated TDDFT for nonequilibrium dynamics 37, a key ingredient of which are the densities of the states involved in the dynamics. In this context, it is of particular interest to explore its performance in fully non-equilibrium dynamics which have a propensity to occupy states of double-excitation character during the evolution, whose densities can be obtained from our formulation.

### v2

@@ 4 Conclusions
In this work we developed a general formalism for obtaining the real-space excited-state densities from linear response TDDFT using time-independent perturbation theory. While the approach is equivalent to earlier work 4, 5 presented using the stationary property of excitation energies and the variational principle, our work here goes beyond the restriction to the adiabatic approximation assumed there. We observe that even with two contrasting approximations (LDA and EXX), we obtain accurate results for the lowest local and charge-transfer excitations between closed-shell fragments. Within the small-matrix approximation, which in its essence retains a single KS transition in its calculation of the energy, the excited-state density still incorporates corrections from all possible KS transitions, leading to a significant improvement over the KS approximation. Our results show the importance of an accurate KS potential in the calculation of excited-state density, which is better provided by adiabatic EXX than the LDA. For the future calculations of excited-state densities in real systems, our results are consistent with previous works 16, 17, 18, 19, 20, 21, 22 that suggest the use of hybrid functionals with large fraction of exact exchange as a more reliable choice.
The practical use of the general formalism applies to the adiabatic approximation and to proposed approximations that are available for frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ kernels. While a general-purpose frequency-dependent kernel does not exist, the formulae here hold for existing frequency-dependent kernels developed for particular situations. We show that for doubly-excited states, the densities from the DSPA and DSMA-based formulation, both previously shown to be good for excitation energies and DSMA for oscillator strengths, provides accurate double excitation densities, which is not achievable through adiabatic approximation. The general idea of obtaining excited state densities from derivatives of energies with respect to the external potential could also be applied to other methods, e.g. pp-RPA 47, 48, BSE 49, 50.
Benchmarking on a wider set of real molecules is required to further probe the efficiency and accuracy of the linear-response formalism developed for real-space excited-state densities for states of both single and double excitation character, as has been explored for dipole moments restricted to the adiabatic approximation (single excitations) in earlier works 16, 17, 18, 19, 20, 21, 22, 23. Beyond static properties, the formalism can be used in the response-reformulated TDDFT for nonequilibrium dynamics 51, a key ingredient of which are the densities of the states involved in the dynamics. In this context, it is of particular interest to explore its performance in fully non-equilibrium dynamics which have a propensity to occupy states of double-excitation character during the evolution, whose densities can be obtained from our formulation.